# On diffusion-based generative models and their error bounds: The log-concave case with full convergence estimates

### v4

@@ C.2 Proof of the Main Result for the General Case
$\begin{split}&\leq\frac{\gamma^{-1}}{2}\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}|b^{(k)}(T-r,\hat{\theta},Y_{r}^{\text{aux}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}})|^{2}\mathrm{d}r\ \mathrm{d}s\right]\\ &\quad+\frac{\gamma}{2}\sum_{k=1}^{M}\int_{0}^{t}\mathbb{E}\left[\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|^{2}\mathrm{d}s\right]\\ &\quad+\sum_{k=1}^{M}\int_{0}^{t}\left[\mathbb{E}\left(\int_{\lfloor s/\gamma\rfloor\gamma}^{s}|b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})|\mathrm{d}r\right)^{2}\right]^{1/2}\\ &\qquad\qquad\quad\times\left[\mathbb{E}\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|^{2}\right]^{1/2}\mathrm{d}s\\ &\quad+\sum_{k=1}^{M}\int_{0}^{t}\left[\mathbb{E}\left(\int_{\lfloor s/\gamma\rfloor\gamma}^{s}2|s^{(k)}(T-r,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})-s^{(k)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})|\ \mathrm{d}r\right)^{2}\right]^{1/2}\\ &\qquad\qquad\quad\times\left[\mathbb{E}\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|^{2}\right]^{1/2}\mathrm{d}s\\ &\leq\mathbb{E}\left[\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\gamma^{-1}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))|Y_{r}^{\text{aux}}-\widehat{Y}_{r}^{\text{EM}}|^{2}\mathrm{d}r\ \mathrm{d}s\right]\\ &\quad+2\gamma^{2}\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\sum_{j=1}^{M}\left|\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\right|^{2}\mathrm{d}s\right]\\ &\quad+\gamma^{1/2}2(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}\\ &\qquad\times\sum_{k=1}^{M}\int_{0}^{t}\left[\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\mathbb{E}[|\widehat{Y}_{r}^{\text{EM}}-\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}}|^{2}]\ \mathrm{d}r\right]^{1/2}\left[\mathbb{E}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\left|\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\right|^{2}\mathrm{d}r\right]^{1/2}\mathrm{d}s\\ &\quad+\gamma^{1+\alpha}4\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}\sum_{k=1}^{M}\int_{0}^{t}\left[\mathbb{E}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\left|\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\right|^{2}\mathrm{d}r\right]^{1/2}\mathrm{d}s\\ &\leq(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))\int_{0}^{t}\sup_{0\leq r\leq s}\mathbb{E}\left[|Y_{r}^{\text{aux}}-\widehat{Y}_{r}^{\text{EM}}|^{2}\right]\mathrm{d}s+4\gamma^{2}tM(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\quad+\gamma^{3/2}2(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}\sum_{k=1}^{M}\int_{0}^{t}\left[\mathbb{E}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\left|\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\right|^{2}\mathrm{d}r\right]^{1/2}\mathrm{d}s\\ &\quad+\gamma^{1+\alpha}4\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}\sum_{k=1}^{M}\int_{0}^{t}\left[\mathbb{E}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\left|\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\right|^{2}\mathrm{d}r\right]^{1/2}\mathrm{d}s\end{split}$
@@
$\begin{split}&\leq(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))\int_{0}^{t}\sup_{0\leq r\leq s}\mathbb{E}\left[|Y_{r}^{\text{aux}}-\widehat{Y}_{r}^{\text{EM}}|^{2}\right]\ \mathrm{d}s+4\gamma^{2}tM(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\quad+[\gamma^{2}2(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+\gamma^{3/2+\alpha}4\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\times[tM\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}].\end{split}$ (71)
Using (68), (69) and (71) in (67), we have
@@
$\begin{split}&\sup_{0\leq s\leq t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\\ &\leq\int_{0}^{t}4(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))\sup_{0\leq r\leq s}\mathbb{E}\left[|Y_{r}^{\text{aux}}-\widehat{Y}_{r}^{\text{EM}}|^{2}\right]\ \text{d}s\\ &\quad+\gamma^{2\alpha}t2\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad\qquad\quad+4\zeta^{-1}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\qquad\qquad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\qquad\qquad\quad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\qquad\qquad\times[M\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}\\ &\leq 2e^{4(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))t}\gamma^{2\alpha}t\\ &\times\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad+4\zeta^{-1}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\quad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\qquad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\quad\times[M\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}.\end{split}$ (72)
Using (72) and $t_{K}=T-\epsilon$, we have
$\begin{split}&W_{2}(\mathscr{L}(Y_{t_{K}}^{\text{aux}}),\mathscr{L}(Y_{K}^{\text{EM}}))\\ &\leq\sqrt{\mathbb{E}\left[|Y_{t_{K}}^{\text{aux}}-\widehat{Y}_{t_{K}}^{\text{EM}}|^{2}\right]}\\ &\leq\sqrt{2}e^{2(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))(T-\epsilon)}\gamma^{\alpha}\sqrt{T-\epsilon}\\ &\quad\times\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad\quad+4\zeta^{-1}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\qquad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\end{split}$
@@
$\begin{split}&\qquad\quad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\qquad\quad\times[M\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}^{1/2}.\end{split}$ (73)
Final upper bound on $W_{2}(\mathscr{L}(Y_{K}^{\text{EM}}),\pi_{\mathsf{D}}).$
Substituting (56), (60), (63), and (73) into (54), we have

### v5

@@ D.2 Proof of the Main Result for the General Case
$\begin{split}&\qquad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(\int_{\lfloor s/\gamma\rfloor\gamma}^{s}(b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}}))\ \mathrm{d}r\right)\\ &\qquad\qquad\qquad\qquad\times\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\ \mathrm{d}s\Bigg{]}\\ &\qquad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\left(\int_{\lfloor s/\gamma\rfloor\gamma}^{s}2(s^{(k)}(T-r,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})-s^{(k)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}}))\ \mathrm{d}r\right)\\ &\qquad\qquad\qquad\qquad\times\left(\sum_{j=1}^{M}\ \frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sqrt{2}\ \mathrm{d}\bar{B}^{(j)}_{r}\right)\ \mathrm{d}s\Bigg{]}\\ &\leq\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\gamma^{-1/2}|b^{(k)}(T-r,\hat{\theta},Y_{r}^{\text{aux}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}})|\mathrm{d}r\\ &\qquad\qquad\qquad\qquad\times\gamma^{1/2}\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|\mathrm{d}s\Bigg{]}\\ &\qquad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}|b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})|\mathrm{d}r\\ &\qquad\qquad\qquad\quad\times\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|\mathrm{d}s\Bigg{]}\\ &\qquad+\sum_{k=1}^{M}\mathbb{E}\Bigg{[}\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}2|s^{(k)}(T-r,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})-s^{(k)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})|\mathrm{d}r\\ &\qquad\qquad\qquad\quad\times\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|\mathrm{d}s\Bigg{]}\\ &\leq\frac{\gamma^{-1}}{2}\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}|b^{(k)}(T-r,\hat{\theta},Y_{r}^{\text{aux}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}})|^{2}\mathrm{d}r\ \mathrm{d}s\right]\\ &\quad+\frac{\gamma}{2}\sum_{k=1}^{M}\int_{0}^{t}\mathbb{E}\left[\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|^{2}\mathrm{d}s\right]\\ &\quad+\sum_{k=1}^{M}\int_{0}^{t}\left[\mathbb{E}\left(\int_{\lfloor s/\gamma\rfloor\gamma}^{s}|b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{r}^{\text{EM}})-b^{(k)}(T-r,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})|\mathrm{d}r\right)^{2}\right]^{1/2}\\ &\qquad\qquad\quad\times\left[\mathbb{E}\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|^{2}\right]^{1/2}\mathrm{d}s\\ &\quad+\sum_{k=1}^{M}\int_{0}^{t}\left[\mathbb{E}\left(\int_{\lfloor s/\gamma\rfloor\gamma}^{s}2|s^{(k)}(T-r,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor r/\gamma\rfloor\gamma})-s^{(k)}(T-\lfloor r/\gamma\rfloor\gamma,\hat{\theta},\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}})|\ \mathrm{d}r\right)^{2}\right]^{1/2}\\ &\qquad\qquad\quad\times\left[\mathbb{E}\left|\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\sqrt{2}\ \mathrm{d}\bar{B}_{r}^{(j)}\right|^{2}\right]^{1/2}\mathrm{d}s\end{split}$
@@
$\begin{split}&\leq\mathbb{E}\left[\int_{0}^{t}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\gamma^{-1}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))|Y_{r}^{\text{aux}}-\widehat{Y}_{r}^{\text{EM}}|^{2}\mathrm{d}r\ \mathrm{d}s\right]\\ &\quad+2\gamma^{2}\sum_{k=1}^{M}\mathbb{E}\left[\int_{0}^{t}\sum_{j=1}^{M}\left|\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\right|^{2}\mathrm{d}s\right]\\ &\quad+\gamma^{1/2}2(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}\\ &\qquad\times\sum_{k=1}^{M}\int_{0}^{t}\left[\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\mathbb{E}[|\widehat{Y}_{r}^{\text{EM}}-\widehat{Y}_{\lfloor r/\gamma\rfloor\gamma}^{\text{EM}}|^{2}]\ \mathrm{d}r\right]^{1/2}\left[\mathbb{E}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\left|\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\right|^{2}\mathrm{d}r\right]^{1/2}\mathrm{d}s\\ &\quad+\gamma^{1+\alpha}4\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}\sum_{k=1}^{M}\int_{0}^{t}\left[\mathbb{E}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\left|\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\right|^{2}\mathrm{d}r\right]^{1/2}\mathrm{d}s\\ &\leq(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))\int_{0}^{t}\sup_{0\leq r\leq s}\mathbb{E}\left[|Y_{r}^{\text{aux}}-\widehat{Y}_{r}^{\text{EM}}|^{2}\right]\mathrm{d}s+4\gamma^{2}tM(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\quad+\gamma^{3/2}2(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}\sum_{k=1}^{M}\int_{0}^{t}\left[\mathbb{E}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\left|\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\right|^{2}\mathrm{d}r\right]^{1/2}\mathrm{d}s\\ &\quad+\gamma^{1+\alpha}4\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}\sum_{k=1}^{M}\int_{0}^{t}\left[\mathbb{E}\int_{\lfloor s/\gamma\rfloor\gamma}^{s}\left|\sum_{j=1}^{M}\frac{\partial b^{(k)}(T-s,\hat{\theta},\widehat{Y}^{\text{EM}}_{\lfloor s/\gamma\rfloor\gamma})}{\partial x^{j}}\right|^{2}\mathrm{d}r\right]^{1/2}\mathrm{d}s\\ &\leq(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))\int_{0}^{t}\sup_{0\leq r\leq s}\mathbb{E}\left[|Y_{r}^{\text{aux}}-\widehat{Y}_{r}^{\text{EM}}|^{2}\right]\ \mathrm{d}s+4\gamma^{2}tM(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\quad+[\gamma^{2}2(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+\gamma^{3/2+\alpha}4\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\times[tM\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}].\end{split}$ (76)
Using 73, 74 and 76 in 72, we have
@@
$\begin{split}&\sup_{0\leq s\leq t}\mathbb{E}\left[|Y_{s}^{\text{aux}}-\widehat{Y}_{s}^{\text{EM}}|^{2}\right]\\ &\leq\int_{0}^{t}4(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))\sup_{0\leq r\leq s}\mathbb{E}\left[|Y_{r}^{\text{aux}}-\widehat{Y}_{r}^{\text{EM}}|^{2}\right]\ \text{d}s\end{split}$
@@
$\ &\quad+\gamma^{2\alpha}t2\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad\qquad\quad+4\zeta^{-1}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\qquad\qquad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\qquad\qquad\quad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\qquad\qquad\times[M\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}\\ &\leq 2e^{4(1+\zeta+\mathsf{K}_{3}(1+2T^{\alpha}+4\mathsf{K}_{3}(1+4T^{2\alpha})))t}\gamma^{2\alpha}t\\ &\times\Bigg{(}\mathsf{K}_{4}^{2}\zeta^{-1}(1+4T^{2\alpha})C_{\mathsf{EMose},4}+4M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))+2\zeta^{-1}\mathsf{K}_{1}^{2}(1+8(\widetilde{\varepsilon}_{\text{AL}}+|\theta^{*}|^{2}))\\ &\qquad+4\zeta^{-1}M(1+8\mathsf{K}^{2}_{3}(1+4T^{2\alpha}))\\ &\qquad\quad\times[(1+16\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha}))C_{\mathsf{EM},2}(T)+32\mathsf{K}^{2}_{\text{Total}}(1+T^{2\alpha})(1+2\widetilde{\varepsilon}_{\text{AL}}+2|\theta^{*}|^{2})]\\ &\qquad+2[(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}C^{1/2}_{\mathsf{EMose},2}+2\mathsf{K}_{1}(1+8\widetilde{\varepsilon}_{\text{AL}}+8|\theta^{*}|^{2})^{1/2}]\\ &\qquad\quad\times[M\sqrt{2}(1+8\mathsf{K}_{3}^{2}(1+4T^{2\alpha}))^{1/2}]\Bigg{)}.\end{split}$ (77)
Using 77 and $t_{K}=T-\epsilon$, we have